OpenStudy (anonymous):

\[f(x)=3(x-1)^{2}-2\] vertex? x-intercept? y-intercept?

OpenStudy (anonymous):

vertex (1,-2)

OpenStudy (anonymous):

x intercepts, set y to 0

OpenStudy (anonymous):

(x-1)^2 = 2/3 x-1= +- sqrt (2/3) x= 1 +-sqrt(2/3)

OpenStudy (anonymous):

y intercepts, let x be zero so y= 3(-1)^2 -2 = 3-2 =1

OpenStudy (anonymous):

The x-intercept has to be a decimal